The contraption below has the following characteristics: m1 = 2.0 x 10^3 kg, m2 = 4.4 x 10^2 kg, the man on top of m2 has a mass m = 6.0 x 10^1 kg, and all surfaces are frictionless. What is the apparent weight of the man?


1
Expert's answer
2021-07-13T11:52:49-0400

Gives

"m_1=2.0\times10^3kg\\m_2=4.4\times10^2kg\\m=6.0\times10^1kg"

"g=9.8m\/sec^2"

We know that

"W=mg"

Put value

"W=6.0\times10^1\times9.8=588N"